how do you write the equation of a line?
(-2,5) slope=-3
equation y-y1=m(x-x1)

Answer by jim_thompson5910(35256)

y-y1=m(x-x1)

y-y1 = -3(x-x1)

y-5 = -3(x-(-2))

y-5 = -3(x+2) <--- equation in point slope form

y-5 = -3x - 6

y = -3x - 6 + 5

y = -3x - 1 <--- equation in slope intercept form